at 7.3% how long to double your money? quadruple it?

Present value Years Interest Rate Future Value
$1 7.3% $2
$1 7.3% $4
Answer #1

t = ln(FV / PV) / ln(1 + r)

t = ln($2 / $1) / ln 1.073 = 9.84 years

t = ln($4 / $1) / ln 1.073 = 19.68 years